If PQ= 2x , QR = 5x , and PR = 7 , what is QR?
1
Expert's answer
2020-09-16T19:26:40-0400

"PQ+QR=PR"

If "PQ=2x", "QR=5x" and "PR=7", then

"2x+5x=7"

"7x=7"

"x=1"

Calculate the value of "QR":

"QR=5x=5*1=5"

Solution: "QR=5".